# Benchmarking Pedestrian Dynamics Models for Common Scenarios: An Evaluation of Force-Based Models

### v1

@@ III Simulation results from force-based models
In this section, the experimental scenarios are reproduced using five force-based models, maintaining the exact initial and final positions observed during the experiments. Data for the same in each scenario are provided at https://github.com/kanika201293/Pedestrian-Experimental-Data. The values of the evaluating parameters are calculated from the simulations to assess the models’ ability to predict the experimental outcomes. Additionally, a parameter for maximum overlap is introduced to provide a more comprehensive evaluation of the models.
@@
Universal Power Law (UPL): The UPL is developed by Karamouzas et al. (2014) karamouzas2014universal , which is formulated after an empirical analysis of two major crowd datasets. The fundamental concept is that the strength of the interaction force is determined only by a single factor, namely, the time-to-collision, denoted as $\tau$. This term inherently influences the angle-dependent nature of interaction force. Note, interaction with walls utilizes the same formula provided that the pedestrian velocity is directed towards the nearest point to the wall, where the nearest point is treated as a static obstacle with a radius equal to $0\;m$.
Social Force Model - circular (SFMc): Helbing is one of the earliest researchers to develop a model to study pedestrian dynamics, i.e., Social Force Model helbing1995social . SFM has been improved occasionally to bring the simulation results closer to reality. SFMc is a straightforward version of the social force model where the interaction force is only a function of surface-to-surface distance between the pedestrian and the entity. The interaction region of the pedestrian in this case is in the form of a circle helbing2000simulating .
Social Force Model - elliptical (SFMe): One of the earliest versions of social force models used the interaction region of a pedestrian in the form of an ellipse, which is oriented towards the direction of motion helbing1995social . In this case, the force is dependent on the distance as well as the velocity of the pedestrian.
The derived formulae and optimized values of parameters for SFMc and SFMe are taken from the article published by Johansson et al. (2007) johansson2007specification .
Centrifugal Force Model - circular (CFMc): The idea of the centrifugal force model was first proposed by Yu et al. (2005) yu2005centrifugal that has a resemblance to the concept of “centrifugal force”. Subsequently, Chraibi expanded upon the model, see chraibi2009quantitative ; chraibi2010generalized . Similar to SFM, CFM also has a circular and elliptical interaction region distinguished by different definitions of the distance between the pedestrian and the obstacle. For CFMc, the radii of circular potential regions of pedestrians are dependent on the respective velocities. Note that in the case of a stationary pedestrian, the radius of the interaction region takes the minimum value of the actual radius of the entity.
@@
Note that, for all the models, the simulation time step size $\Delta t$ is set as $0.01\;s$, the maximum time step size that avoids numerical instability (refer to Appendix section A). This study assigns specific colors for each force-based model and the experimental data. The color coding remains consistent throughout the article.
@@
Table 3: Comparison of the five force-based models with experimental data using different evaluating parameters for SOSP
@@
However, in real-life scenarios, it is unlikely that a person stands precisely in the expected position with no offset. Trajectories and normalized speed profiles for one of the data are plotted in Figure 4 (multimedia available online). In this case, the volunteer started with an offset as low as 11 mm, yet there were no sudden changes in the trajectory or speed during the experiment, unlike simulation results. As explained next, all models showed significant slowdown of the pedestrian as it approached the obstacle in the simulations, which is in sharp contrast to the experimental observations.
While simulating the experimental scenarios using force-based models using the same initial and final positions, we calculated and reported the corresponding evaluating parameters in Table 3. Model SFMc predicted over 80% unsuccessful trajectories, SFMe showed less than 2% unsuccessful trajectories, while all other models had all successful trajectories. The large number of unsuccessful trajectories for model SFMc is primarily due to high overlap, with a maximum overlap of 100%. This high failure rate stems from SFMc’s weak interaction force, leading to an inability to avoid obstacles effectively, as depicted in Figure 4(a) (multimedia available online). Some overlap was observed in models SFMe and CFMc, but no overlap occurred with UPL and CFMe. No backward motion or oscillation was reported for any model. However, sudden changes in direction were noted under the ‘path smoothness’ parameter, with model SFMe showing a maximum change exceeding 90 degrees, as illustrated in the trajectory. It is important to note that human walking is inherently less smooth than simulated motion, which can result in lower path smoothness values in simulations compared to experimental values in some cases. Large speed deviations from the desired speed were observed in models SFMe, CFMc, and CFMe, evident in the velocity dips highlighted in Figure 4(b). These speed slowdowns significantly increased travel time for these three models, and respective values are provided in Table 3.
Table 4: Comparison of the five force-based models with experimental data using different evaluating parameters for Head-On.
@@
Head-On: One example from the head-on experiments is plotted in Figure 5 (multimedia available online). In trajectory plots, solid lines with filled circles represent pedestrian A, and dashed lines with empty circles denote pedestrian B. Table 4 reports the evaluating parameters for head-on scenarios simulated using force-based models. Unsuccessful cases are predicted for models SFMc, SFMe, and CFMe, with high overlapping proportions. For SFMc and SFMe, the overlap is also evident from the trajectories shown in Figure 5(a) (multimedia available online). Models UPL and CFMc report 0% and less than 1% maximum overlap, respectively, and no oscillation is observed in any model. Large sudden deviations in the direction of motion are observed in models SFMe and CFMe, also shown in the trajectory. The parameter of path smoothness measures the sudden change in the direction of motion rather than the overall change in the path. Hence, the value of path smoothness is lower for CFMc relative to CFMe. Although the overall deviation is large for CFMc, it is smoother compared to the abrupt changes observed in CFMe’s trajectory at around $x=4\;m$ in Figure 5(a) (multimedia available online). Significant speed deviations are reported for models CFMc and CFMe, as can also be seen in the velocity profiles of Figure 5(b). Consequently, these models also exhibit longer travel times, as indicated in Table 4. For other models, the simulation predicts travel times similar to the experiments.
Table 5: Comparison of the five force-based models with experimental data using different evaluating parameters for Parallel-Ped.
@@
Table 6: Comparison of the five force-based models with experimental data using different evaluating parameters for MOSP.
@@
Parallel-Ped: In this scenario, two pedestrians, A and B, share a common goal, as illustrated in Figure 2(c). One example of speed profiles and trajectories for both pedestrians generated using force-based models are provided in Figure 6 (multimedia available online). In trajectory plots, solid lines with filled circles represent pedestrian A, while dashed lines with empty circles represent pedestrian B. Evaluating parameters obtained from simulations are shown in Table 5. The model SFMc predicted more than 30% unsuccessful trajectories due to large overlap in most cases, with a maximum overlap of 95.25%. For scenarios SOSP, head-on, and parallel ped, involving only two entities, the main cause of unsuccessful trajectories is overlap rather than the pedestrian getting stuck. Oscillations are not predicted in any model for this scenario. The trajectories of pedestrian A in Figure 6(a) (multimedia available online) indicate that, for most models, the faster pedestrian A is unable to overtake the slower pedestrian B, resulting in mostly straight-line trajectories. However, in experiments, pedestrian A overtakes B, in contrast to these simulated predictions. Thus, path smoothness for most models is better than the experimental value, except for UPL. In the UPL model, pedestrian A successfully overtakes pedestrian B. However, pedestrian B also changes the path direction to allow passage for pedestrian A, unlike during the experiment where A finds its own path without affecting B’s motion. This is logical as pedestrian B, positioned ahead of A, cannot see A and should not alter its path. For UPL, sudden direction change for both pedestrians ranges from 2 to 4 degrees. Models SFMe, CFMc, and CFMe exhibit significant speed deviations from the desired speed, with maximum deviations between 20% and 48%. Large deviation in velocities from the desired speed ($V_{des}$) is also observed in the normalized speed plots given in Figure 6(b)-(c). For these models, pedestrian A cannot overtake B, resulting in a slower speed of A (Figure 6(b)), while the speed of B is maintained for models CFMc and CFMe (Figure 6(c)). The SFMe model shows an unrealistic push effect on pedestrian B, increasing the speed by up to 1.4 times. The models predicted travel times similar to the experiment.
@@
MOSP: In the MOSP scenario, a single pedestrian passes through a maze of randomly placed obstacles to reach the goal, with experiments conducted across four area fractions: 3.74% (Case A), 6.54% (Case B), 11.22% (Case C), and 14.96% (Case D). Evaluating parameters calculated by reproducing each setup for the exact initial and final conditions are reported in Table 6. These show that as the area fraction covered by the obstacles increases, the number of unsuccessful trajectories also increases. (i.e. the model predictions become worse with increasing density of stationary obstacles). Note, such situations of pedestrians walking around several other stationary obstacles (or other standing pedestrians) are extremely common across crowded public spaces, like railway stations, bus stations, food joints etc. In MOSP, the large number of obstacles leads to unsuccessful trajectories due to a combination of three factors mentioned in Section II.2. For the SFMe model, no successful trajectories were observed in Cases C and D, as high repulsive forces prevented pedestrians from entering the measurement region, resulting in ‘Not Attempted (NA)’ for these cases, (refer to Appendix Figure B-5(b)). Surprisingly, the SFMc model showed an increase of successful trajectories from Case C to D due to the obstacle configuration favorable for the model, allowing the pedestrian to reach the goal with low overlap, even when there is a minimal deviation in path due to the weak repulsive forces of the model (refer to Appendix Figure B-5(a)). Both average and maximum overlaps increased with area fraction. However, this parameter is influenced by obstacle configuration, pedestrian position, and the number of interacting obstacles. Thus, this trend should not be seen as a rule. Insignificant oscillation or backward motion was observed across all models. As the number of obstacles increased, the number of turns required to pass through the maze also increased, leading to higher sudden changes in motion direction, reflected in large path smoothness values. Due to frequent changes in path direction, speed deviation from the desired speed also increases with the area fraction. Normalized speed graphs for the four cases, averaged over all the trajectories, can be observed in Appendix Figure B-4. However, the SFMc model reported low path smoothness and speed deviation, resulting in increased travel time with area fraction for all models except SFMc. However, note that the proportion of successful trajectories with SFMc is pretty low ($<25\%$) for both area fractions.
@@
IV Evaluation of Force-Based Models
@@
[…]
@@
Parameters $\boldsymbol{\rightarrow}$ Avg. Overlap Avg. Oscillation Path Smoothness Speed Deviation Travel Time
Exp. Setup $\boldsymbol{\downarrow}$ Cutoff UPL CFMc CFMe Cutoff UPL CFMc CFMe Cutoff UPL CFMc CFMe Cutoff UPL CFMc CFMe Cutoff UPL CFMc CFMe
SOSP $<10\%$checkmark $\checkmark$ $\checkmark$ $<0.005$checkmark $\checkmark$ $\checkmark$ $<1.73^{\circ}$$\boldsymbol{\times}$checkmark $\checkmark$ $<0.14$checkmark $\boldsymbol{\times}$$\boldsymbol{\times}$ $<7.19s$checkmark $\checkmark$ $\checkmark$
Head-On $<10\%$checkmark $\checkmark$ $\checkmark$ $<0.005$checkmark $\checkmark$ $\checkmark$ $<0.91^{\circ}$$\boldsymbol{\times}$checkmark $\boldsymbol{\times}$ $<0.05$checkmark $\boldsymbol{\times}$$\boldsymbol{\times}$ $<4.84s$checkmark $\boldsymbol{\times}$$\boldsymbol{\times}$
Parallel-Ped $<10\%$checkmark $\checkmark$ $\checkmark$ $<0.005$checkmark $\checkmark$ $\checkmark$ $<1.31^{\circ}$$\boldsymbol{\times}$checkmark $\checkmark$ $<0.17$checkmark $\boldsymbol{\times}$$\boldsymbol{\times}$ $<9.41s$checkmark $\checkmark$ $\checkmark$
@@
Case A $<10\%$checkmark $\checkmark$ $\checkmark$ $<0.005$ $\boldsymbol{\times}$checkmark $\checkmark$ $<1.29^{\circ}$$\boldsymbol{\times}$checkmark $\checkmark$ $<0.14$checkmark $\boldsymbol{\times}$$\boldsymbol{\times}$ $<5.81s$checkmark $\boldsymbol{\times}$$\boldsymbol{\times}$
Case B $<10\%$checkmark $\checkmark$ $\checkmark$ $<0.005$ $\boldsymbol{\times}$checkmark $\checkmark$ $<2.1^{\circ}$$\boldsymbol{\times}$$\boldsymbol{\times}$$\boldsymbol{\times}$ $<0.35$checkmark $\boldsymbol{\times}$$\boldsymbol{\times}$ $<5.96s$checkmark $\boldsymbol{\times}$$\boldsymbol{\times}$
Case C $<10\%$checkmark $\checkmark$ $\boldsymbol{\times}$ $<0.005$ $\boldsymbol{\times}$checkmark $\checkmark$ $<2.61^{\circ}$$\boldsymbol{\times}$$\boldsymbol{\times}$$\boldsymbol{\times}$ $<0.37$ $\boldsymbol{\times}$$\boldsymbol{\times}$$\boldsymbol{\times}$ $<6.74s$$\boldsymbol{\times}$$\boldsymbol{\times}$$\boldsymbol{\times}$
Case D $<10\%$checkmark $\boldsymbol{\times}$$\boldsymbol{\times}$ $<0.005$ $\boldsymbol{\times}$checkmark $\checkmark$ $<2.01^{\circ}$$\boldsymbol{\times}$$\boldsymbol{\times}$$\boldsymbol{\times}$ $<0.24$ $\boldsymbol{\times}$$\boldsymbol{\times}$$\boldsymbol{\times}$ $<6.53s$$\boldsymbol{\times}$$\boldsymbol{\times}$$\boldsymbol{\times}$
@@
[…]
In the second stage, $\checkmark$ and $\times$ marks are again assigned to the shortlisted models, based on whether their average parameter values fall within the specified cutoff values (refer to Table 8). Cutoff values for path smoothness, speed deviation, and travel time are determined by using the $3\sigma$ rule, adding three standard deviations to the mean parameter values observed during the experiments, while the cutoff values for overlap and oscillation parameters are set at 10% and 0.005, respectively, due to no observed values for these parameters during experiments.
Normalized success score for a parameter is derived by counting the number of $\checkmark$ it receives (refer to Table 9). The results after the two-stage process show that UPL yields the best outcomes among the five force-based models, followed by CFMc and CFMe. However, the final scores of UPL, CFMc and CFMe are separated by $<10\%$, suggesting that they yield comparable predictions for the situations considered here. It is important to recognize the significance of the first stage of shortlisting. If scoring had been directly applied to all models, SFMc would have received the highest score despite failing to achieve 80% successful trajectories for all the setups, including the simplest SOSP scenario. However, our analysis also indicates the failure of all models for the higher-density MOSP scenarios.

### v2

@@ III Simulation results from force-based models
In this section, the experimental scenarios are reproduced using five force-based models, maintaining the exact initial and final positions observed during the experiments. Data for the same in each scenario are provided at https://github.com/kanika201293/Pedestrian-Experimental-Data. The values of the evaluating metrics are calculated from the simulations to assess the models’ ability to predict the experimental outcomes. Additionally, a metric for maximum overlap is introduced to provide a more comprehensive evaluation of the models.
@@
Universal Power Law (UPL): The UPL is developed by Karamouzas et al. (2014) Karamouzas et al. (2014), which is formulated after an empirical analysis of two major crowd datasets. The fundamental concept is that the strength of the interaction force is determined only by a single factor, namely, the time-to-collision, denoted as $\tau$. This term inherently influences the angle-dependent nature of interaction force. Note, interaction with walls utilizes the same formula provided that the pedestrian velocity is directed towards the nearest point to the wall, where the nearest point is treated as a static obstacle with a radius equal to 0m0\;m.
Social Force Model - circular (SFMc): Helbing is one of the earliest researchers to develop a model to study pedestrian dynamics, i.e., Social Force Model Helbing and Molnar (1995). SFM has been improved occasionally to bring the simulation results closer to reality. SFMc is a straightforward version of the social force model where the interaction force is only a function of surface-to-surface distance between the pedestrian and the entity. The interaction region of the pedestrian in this case is in the form of a circle Helbing et al. (2000).
Social Force Model - elliptical (SFMe): One of the earliest versions of social force models used the interaction region of a pedestrian in the form of an ellipse, which is oriented towards the direction of motion Helbing and Molnar (1995). In this case, the force is dependent on the distance as well as the velocity of the pedestrian.
The derived formulae and optimized values of parameters for SFMc and SFMe are taken from the article published by Johansson et al. (2007) Johansson et al. (2007).
Centrifugal Force Model - circular (CFMc): The idea of the centrifugal force model was first proposed by Yu et al. (2005) Yu et al. (2005) that has a resemblance to the concept of “centrifugal force”. Subsequently, Chraibi expanded upon the model, see Chraibi et al. (2009, 2010). Similar to SFM, CFM also has a circular and elliptical interaction region distinguished by different definitions of the distance between the pedestrian and the obstacle. For CFMc, the radii of circular potential regions of pedestrians are dependent on the respective velocities. Note that in the case of a stationary pedestrian, the radius of the interaction region takes the minimum value of the actual radius of the entity.
@@
Note that, for all the models, the simulation time step size $\Delta t$ is set as 0.01s0.01\;s, the maximum time step size that avoids numerical instability (refer to Supplementary material). This study assigns specific colors for each force-based model and the experimental data. The color coding remains consistent throughout the article.
@@
Table 3: Comparison of the five force-based models with experimental data using different evaluating metrics for SOSP
@@
However, in real-life scenarios, it is unlikely that a person stands precisely in the expected position with no offset. Trajectories and normalized speed profiles for one of the data are plotted in Figure 4 (multimedia available online). In this case, the volunteer in the experiment started with an offset as low as 11 mm, yet there were no sudden changes in the trajectory or speed during the experiment, unlike simulation results. As explained next, all models showed significant slowdown of the pedestrian as it approached the obstacle in the simulations, which is in sharp contrast to the experimental observations. The possible reason behind such behavior is force-based models relying on forces to navigate around the obstacles, which naturally leads to a change in the magnitude of the velocity (speed), as can be seen in Figure 4(b).
While simulating the experimental scenarios using force-based models using the same initial and final positions, we calculated and reported the corresponding evaluating metrics in Table 3. Model SFMc predicted over 80% unsuccessful trajectories, SFMe showed less than 2% unsuccessful trajectories, while all other models had all successful trajectories. The large number of unsuccessful trajectories for model SFMc is primarily due to high overlap, with a maximum overlap of 100%. This high failure rate stems from SFMc’s weak interaction force, leading to an inability to avoid obstacles effectively, as depicted in Figure 4(a) (multimedia available online). Some overlap was observed in models SFMe and CFMc, but no overlap occurred with UPL and CFMe. No backward motion or oscillation was reported for any model. However, sudden changes in direction were noted under the ‘path smoothness’ metric, with model SFMe showing a maximum change exceeding 90 degrees, as illustrated in the trajectory. It is important to note that human walking is inherently less smooth than simulated motion, which can result in lower path smoothness values in simulations compared to experimental values in some cases. Large speed deviations from the desired speed were observed in models SFMe, CFMc, and CFMe, evident in the velocity dips highlighted in Figure 4(b). These speed slowdowns significantly increased travel time for these three models, and respective values are provided in Table 3.
Table 4: Comparison of the five force-based models with experimental data using different evaluating metrics for Head-On.
@@
Head-On: One example from the head-on experiments is plotted in Figure 5 (multimedia available online). In trajectory plots, solid lines with filled circles represent pedestrian A, and dashed lines with empty circles denote pedestrian B. Table 4 reports the evaluating metrics for head-on scenarios simulated using force-based models. Unsuccessful cases are predicted for models SFMc, SFMe, and CFMe, with high overlapping proportions. For SFMc and SFMe, the overlap is also evident from the trajectories shown in Figure 5(a) (multimedia available online). Models UPL and CFMc report 0% and less than 1% maximum overlap, respectively, and no oscillation is observed in any model. Large sudden deviations in the direction of motion are observed in models SFMe and CFMe, also shown in the trajectory. The metric of path smoothness measures the sudden change in the direction of motion rather than the overall change in the path. Hence, the value of path smoothness is lower for CFMc relative to CFMe. Although the overall deviation is large for CFMc, it is smoother compared to the abrupt changes observed in CFMe’s trajectory at around x=4mx=4\;m in Figure 5(a) (multimedia available online). Significant speed deviations are reported for models CFMc and CFMe, as can also be seen in the velocity profiles of Figure 5(b). Consequently, these models also exhibit longer travel times, as indicated in Table 4. For other models, the simulation predicts travel times similar to the experiments.
Table 5: Comparison of the five force-based models with experimental data using different evaluating metrics for Parallel-Ped.
@@
Table 6: Comparison of the five force-based models with experimental data using different evaluating metrics for MOSP.
@@
Parallel-Ped: In this scenario, two pedestrians, A and B, share a common goal, as illustrated in Figure 2(c). One example of speed profiles and trajectories for both pedestrians generated using force-based models are provided in Figure 6 (multimedia available online). In trajectory plots, solid lines with filled circles represent pedestrian A, while dashed lines with empty circles represent pedestrian B. Evaluating metrics obtained from simulations are shown in Table 5. The model SFMc predicted more than 30% unsuccessful trajectories due to large overlap in most cases, with a maximum overlap of 95.25%. For scenarios SOSP, head-on, and parallel ped, involving only two entities, the main cause of unsuccessful trajectories is overlap rather than the pedestrian getting stuck. Oscillations are not predicted in any model for this scenario. The trajectories of pedestrian A in Figure 6(a) (multimedia available online) indicate that, for most models, the faster pedestrian A is unable to overtake the slower pedestrian B, resulting in mostly straight-line trajectories. However, in experiments, pedestrian A overtakes B, in contrast to these simulated predictions. Thus, path smoothness for most models is better than the experimental value, except for UPL. In the UPL model, pedestrian A successfully overtakes pedestrian B. However, pedestrian B also changes the path direction to allow passage for pedestrian A, unlike during the experiment where A finds its own path without affecting B’s motion. This is logical as pedestrian B, positioned ahead of A, cannot see A and should not alter its path. For UPL, sudden direction change for both pedestrians ranges from 2 to 4 degrees. Models SFMe, CFMc, and CFMe exhibit significant speed deviations from the desired speed, with maximum deviations between 20% and 48%. Large deviation in velocities from the desired speed ($V_{des}$) is also observed in the normalized speed plots given in Figure 6(b)-(c). For these models, pedestrian A cannot overtake B, resulting in a slower speed of A (Figure 6(b)), while the speed of B is maintained for models CFMc and CFMe (Figure 6(c)). The SFMe model shows an unrealistic push effect on pedestrian B, increasing the speed by up to 1.4 times. The models predicted travel times similar to the experiment.
@@
MOSP: In the MOSP scenario, a single pedestrian passes through a maze of randomly placed obstacles to reach the goal, with experiments conducted across four area fractions: 3.74% (Case A), 6.54% (Case B), 11.22% (Case C), and 14.96% (Case D). Evaluating metrics calculated by reproducing each setup for the exact initial and final conditions are reported in Table 6. These show that as the area fraction covered by the obstacles increases, the number of unsuccessful trajectories also increases. While it is commonly reported in the literature that force-based models perform well in complex scenarios such as lane formation under moderate to high densities, our findings indicate a decline in model performance as density increases. However, one difference in our situations considered here is the aspect of navigation around a crowded zone with static obstacles, unlike the dynamics in a crowd where every pedestrian is mobile.
Note, such situations of pedestrians walking around several other stationary obstacles (or other standing pedestrians) are extremely common across crowded public spaces, like railway stations, bus stations, food joints etc. In MOSP, the large number of obstacles leads to unsuccessful trajectories due to a combination of three factors mentioned in Section II.2. For the SFMe model, no successful trajectories were observed in Cases C and D, as high repulsive forces prevented pedestrians from entering the measurement region, resulting in ‘Not Attempted (NA)’ for these cases, (refer to Appendix Figure B-5(b)). Surprisingly, the SFMc model showed an increase of successful trajectories from Case C to D due to the obstacle configuration favorable for the model, allowing the pedestrian to reach the goal with low overlap, even when there is a minimal deviation in path due to the weak repulsive forces of the model (refer to Supplementary material Figure B-5(a)). Both average and maximum overlaps increased with area fraction. However, this metric is influenced by obstacle configuration, pedestrian position, and the number of interacting obstacles. Thus, this trend should not be seen as a rule. Insignificant oscillation or backward motion was observed across all models. As the number of obstacles increased, the number of turns required to pass through the maze also increased, leading to higher sudden changes in motion direction, reflected in large path smoothness values. Due to frequent changes in path direction, speed deviation from the desired speed also increases with the area fraction. Normalized speed graphs for the four cases, averaged over all the trajectories, can be observed in Supplementary material Figure B-4. However, the SFMc model reported low path smoothness and speed deviation, resulting in increased travel time with area fraction for all models except SFMc. However, note that the proportion of successful trajectories with SFMc is pretty low ($25\%$) for both area fractions.
@@
IV Evaluation of Models
@@
  • Stage I: The models must meet an eligibility criterion of at least 80% successful trajectories (Table 7).
  • Stage II: The shortlisted models are then evaluated using a scoring system, where a model’s score is determined by the number of times its evaluating metrics fall within the specified cutoff values (Table 8, 9).
Table 7: Stage I: Models meeting eligibility criteria are reported in this table. For each experimental setup, models achieving more than 80% successful cases are marked as $\checkmark$, otherwise ×. Models with the number of $\checkmark$ greater than or equal to the number of × are shortlisted for further evaluation in Stage II, given in Table 8.
@@
Table 8: Stage II: Labels $\checkmark$ and × are assigned to the shortlisted models from Stage I, based on whether their average evaluating metric values fall within the specified cutoff values or not, respectively. The cutoff values are mentioned in the first column corresponding to each metric. The values for overlap and oscillation metrics are set as 10% and 0.005, respectively, for each experimental setup. The cutoff values are determined using the $3\sigma$ rule for the other three metrics, adding three standard deviations to the mean values observed during the experiments. The final scoring for the models is provided in Table 9
@@
[…]
@@
[…]
Table 9: A model’s score for final evaluation is calculated by counting the number of $\checkmark$ it receives in Table 8. The evaluation process highlights that UPL shows the best results among the five force-based models, followed by CFMc and CFMe, while SFMc and SFMe are ‘Not Eligible (NE)’ from Stage I.
@@
In Table 7, the first stage of the models’ eligibility criteria is reported. For each experimental scenario, models achieving more than 80% successful cases are marked as $\checkmark$; otherwise, they are marked as $\times$. Models with a number of $\checkmark$ greater than or equal to the number of $\times$ are shortlisted for further evaluation in Stage II; Otherwise considered as ‘Not Eligible (NE)’. This eligibility criterion primarily determines if a model can provide simulation results comparable to the experimental results for the situations considered. Table 7 clearly indicates that the UPL, CFMc, and CFMe models perform well ($80\%$ successful trajectories) for most situations considered. Note that this study does not imply that the remaining models are entirely ineffective, but rather that they are not suitable for these common situations with their current parametric values (refer to Section C in the Supplementary material for more details). A search across the parameter space of the models is beyond the scope of this article. Interestingly, almost all models are unsuccessful for the MOSP scenarios with higher obstacle densities.
In the second stage, $\checkmark$ and $\times$ marks are again assigned to the shortlisted models, based on whether their average metric values fall within the specified cutoff values (refer to Table 8). Cutoff values for path smoothness, speed deviation, and travel time are determined by using the $3\sigma$ rule, adding three standard deviations to the mean values observed during the experiments, while the cutoff values for overlap and oscillation metrics are set at 10% and 0.005, respectively, due to no observed values for these metrics during experiments.
Normalized success score for a model is derived by counting the number of $\checkmark$ it receives (refer to Table 9). The results after the two-stage process show that UPL yields the best outcomes among the five force-based models, followed by CFMc and CFMe. However, the final scores of UPL, CFMc and CFMe are separated by $10\%$, suggesting that they yield comparable predictions for the situations considered here. It is important to recognize the significance of the first stage of shortlisting. If scoring had been directly applied to all models, SFMc would have received the highest score despite failing to achieve 80% successful trajectories for all the setups, including the simplest SOSP scenario. However, our analysis also indicates the failure of all models for the higher-density MOSP scenarios.